# TBGRecall: A Generative Retrieval Model for E-commerce Recommendation Scenarios

### v1

@@ 1. Introduction
Recommendation systems (Isinkaye et al., 2015; Ko et al., 2022; Gao et al., 2023) are critical to e-commerce (Pfadler et al., 2020) and social platforms (Fan et al., 2019). The retrieval stage (Huang et al., 2024), the pipeline’s foundational step, generates a large candidate item set (e.g., thousands) for subsequent ranking modules, including rough ranking and hybrid ranking. Notably, this process resembles a generative task, making it amenable to modern generative modeling paradigms.
Generative neural models, which excel in NLP (Chen et al., [n. d.]; Touvron et al., 2023) and CV (Radford et al., 2021; Liang et al., 2024), have recently gained traction in recommendation systems. A prominent direction involves leveraging large language models (LLMs): some works use LLMs for feature enhancement in retrieval tasks (Lee et al., 2025; Zhang et al., 2024), while others explore direct LLM-based item generation (Deng et al., 2025). However, challenges like semantic alignment limit their efficacy. Alternative approaches adapt LLM paradigms for recommendations, such as Tiger’s(Rajput et al., 2023) RQ-VAE-based semantic ID framework, later extended in (Zheng et al., 2024; Yang et al., 2025). Other studies propose sparse ID generation via models like HSTU (Zhai et al., 2024) and E4SRec (Li et al., 2023).
@@
To address this, we propose Taobao Generative Recall (TBGRecall), a session-wise autoregressive model where item generation is independent. Its core innovation is Next Session Prediction (NSP), a generation paradigm optimized for retrieval. By reengineering sample structure and loss functions, NSP outperforms classical Next Token Prediction (NTP). Built on HSTU (Zhai et al., 2024), our framework partitions user sequences into multi-session segments, each starting with a session token followed by item tokens. During training, contrastive loss between context tokens and positive/negative item samples replaces regressive loss. At inference, the session token guides Approximate Nearest Neighbor (ANN) search for retrieval. This design resolves autoregressive dependency conflicts and aligns with retrieval infrastructure.
@@ 2.1. Traditional Retrieval Methods
In recommendation systems, the retrieval phase serves as a critical first step to filter vast item corpora into a relevant subset for users, based on historical interactions and behavioral patterns. Traditional approaches include content-based filtering—which leverages item attributes (e.g., keywords, categories) to recommend similar items—and collaborative filtering, which exploits user-item interaction data to identify preferences through community behavior (Sarwar et al., 2001). Notably, item-based collaborative filtering (Sarwar et al., 2001) and matrix factorization (MF) (Koren et al., 2009) demonstrate superior performance by modeling latent relationships and integrating implicit feedback, temporal dynamics, and confidence signals. Large-scale applications further advance these ideas: the Swing algorithm (Yang et al., 2020) constructs product graphs to capture item substitutability and complementarity in e-commerce environments like Taobao. Recent breakthroughs, however, are driven by deep learning-based embedding models, particularly two-tower architectures deployed at YouTube (Covington et al., 2016; Yi et al., 2019), which enhance candidate generation through unbiased representation learning from streaming data while addressing sampling biases. These advancements highlight the transformative impact of deep learning on scalability and precision in large-scale recommendation systems.
@@ 2.2. Generative Models for Recommendation
Recent advances in LLMs enable semantic feature augmentation and generative modeling for recommendation systems. Google proposes a training-free framework integrating LLM embeddings via two-stage retrieval-ranking(Lee et al., 2025). Meta’s EmbSum(Zhang et al., 2024) uses a pretrained encoder-decoder with poly-attention to generate UPE/CPE, refined by LLM-supervised summarization. Bytedance’s HLLM(Chen et al., 2024) decouples content/behavior analysis via Item/User LLMs, improving sequential modeling accuracy.
In sequential recommendation, generative models—particularly sequence-to-sequence architectures—have emerged as a transformative paradigm by framing next-item prediction as autoregressive generation. Early Transformer-based approaches like SASRec demonstrated the efficacy of self-attention in modeling masked user sequences. Building on this, google(Rajput et al., 2023) introduces a content-quantized framework encoding item semantics into discrete semantic-ID tuples, with a decoder generating these tokens to enhance cold-start robustness and diversity control. Tencent(Zheng et al., 2024) bridges language and collaborative signals via vector quantization and alignment tuning, better integrating recommendation-specific features into LLMs. Kuaishou(Deng et al., 2025) further unifies retrieval and ranking within a single model through sparse MoE, session-wise generation, and DPO-based alignment. Recent works extend these ideas: Meta(Zhai et al., 2024) treats recommendation as sequential transduction, enabling efficient scaling in high-cardinality environments(Zhang et al., 2023), while others(Li et al., 2023) combines sequential pretraining with LLM instruction tuning by embedding item IDs into prompts, enabling both accurate retrieval and interpretable recommendations.
@@ 3.3. Next Session Prediction Framework
Given TBGRecall’s alignment with mainstream generative models, we integrate two established LLM techniques—Multi-Token Prediction (MTP) and Mixture-of-Experts (MoE)—to enhance performance. We implement Multi-Session Prediction (MSP) along the session dimension for context tokens, mirroring the MTP paradigm (DeepSeek-AI, 2024; Gloeckle et al., 2024). This introduces extended training signals, enabling explicit modeling of long-range user behavior dependencies and transitional relationships across distant contextual scenarios. For the feed-forward module, we adopt DeepSeekMoE with auxiliary-loss-free load balancing, which harnesses specialized expert networks to acquire richer, task-specific knowledge at comparable inference FLOPS, thereby improving overall model efficacy.
@@ 5.1.1. Dataset
We conduct experiments on two large-scale datasets: 1) RecFlow (Liu et al., 2024), which contains rich contextual information from the Kuaishou. 2) TaoBao dataset, a real-world, industrial-scale dataset derived from traffic logs of TaoBao. Table 1 summarizes the main statistics for these two datasets.
@@ 5.1.3. Baselines
To validate the superior performance of our model, we selected multiple generative recommendation baseline models. The models we selected are all non-LLM architectures that, like TBGRecall, utilize sparse IDs. HSTU, SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019) are sequential modeling-based recommendation models, while YouTubeDNN (Covington et al., 2016) and FDSA (Zhang et al., 2019) represent traditional recommendation approaches. The ONLINE model adopts Taobao’s production-level dual-tower baseline for recommendation retrieval. The ONLINE (DT) denotes that the model is updated through daily training over multiple years, ensuring sustained adaptation to long-term trends and dynamic user preferences.
@@ 5.3. Ablation Study
We conducted comprehensive ablation experiments on key components including TSN, MSP, MoE(DeepSeek-AI, 2024), and RoPE(Su et al., 2024) to assess their individual contributions to model performance. Results confirm that each module independently enhances retrieval accuracy, with the full framework achieving the highest overall effectiveness. MSP contributes the most significant performance improvement, primarily because it increases the amount of information encoded within a single session, enabling the model to better understand the contextual semantics and avoid falling into local decision-making patterns during training. Both session-wise RoPE (Rotary Position Embedding) and MoE (Mixture of Experts) are widely adopted techniques in current state-of-the-art large language models. By incorporating them into our architecture, we achieve considerable gains in model performance. Additionally, the TSN (Token-Specific Network) module enables context tokens and item tokens to align more effectively and efficiently, further enhancing the overall performance of the model. These findings validate the architectural soundness and practical value of the proposed components in advancing generative retrieval systems for e-commerce recommendation.

### v2

@@ 1. Introduction
Recommendation systems (isinkaye2015recommendation; ko2022survey; gao2023survey) are critical to e-commerce (pfadler2020billion) and social platforms (fan2019graph). The retrieval stage (huang2024comprehensive), the pipeline’s foundational step, generates a large candidate item set (e.g., thousands) for subsequent ranking modules, including rough ranking and hybrid ranking. Notably, this process resembles a generative task, making it amenable to modern generative modeling paradigms.
Generative neural models, which excel in NLP (chenlonglora; touvron2023llama) and CV (radford2021learning; liang2024survey), have recently gained traction in recommendation systems. A prominent direction involves leveraging large language models (LLMs): some works use LLMs for feature enhancement in retrieval tasks (googleStar; metaEmbsum), while others explore direct LLM-based item generation (kuaishouOnerec). However, challenges like semantic alignment limit their efficacy. Alternative approaches adapt LLM paradigms for recommendations, such as Tiger’s(googleTiger) RQ-VAE-based semantic ID framework, later extended in (tencentLCRec; baiduCOBRA). Other studies propose sparse ID generation via models like HSTU (metaHSTU) and E4SRec (thuE4SRec). And some other research like sessionrec (huang2025sessionrec) uses new methods for generative recommend.
@@
To address this, we propose Taobao Generative Recall (TBGRecall), a session-wise autoregressive model where item generation is independent. Its core innovation is Next Session Prediction (NSP), a generation paradigm optimized for retrieval. By reengineering sample structure and loss functions, NSP outperforms classical Next Token Prediction (NTP). Built on HSTU (metaHSTU), our framework partitions user sequences into multi-session segments, each starting with a session token followed by item tokens. During training, contrastive loss between context tokens and positive/negative item samples replaces regressive loss. At inference, the session token guides Approximate Nearest Neighbor (ANN) search for retrieval. This design resolves autoregressive dependency conflicts and aligns with retrieval infrastructure.
@@ 2.1. Traditional Retrieval Methods
In recommendation systems, the retrieval phase serves as a critical first step to filter vast item corpora into a relevant subset for users, based on historical interactions and behavioral patterns. Traditional approaches include content-based filtering—which leverages item attributes (e.g., keywords, categories) to recommend similar items—and collaborative filtering, which exploits user-item interaction data to identify preferences through community behavior (CF). Notably, item-based collaborative filtering (CF) and matrix factorization (MF) (MF) demonstrate superior performance by modeling latent relationships and integrating implicit feedback, temporal dynamics, and confidence signals. Large-scale applications further advance these ideas: the Swing algorithm (SWING) constructs product graphs to capture item substitutability and complementarity in e-commerce environments like Taobao. Recent breakthroughs, however, are driven by deep learning-based embedding models, particularly two-tower architectures deployed at YouTube (youtubednn; TTM2), which enhance candidate generation through unbiased representation learning from streaming data while addressing sampling biases. These advancements highlight the transformative impact of deep learning on scalability and precision in large-scale recommendation systems.
@@ 2.2. Generative Models for Recommendation
Recent advances in LLMs enable semantic feature augmentation and generative modeling for recommendation systems. Google proposes a training-free framework integrating LLM embeddings via two-stage retrieval-ranking(googleStar). Meta’s EmbSum(metaEmbsum) uses a pretrained encoder-decoder with poly-attention to generate UPE/CPE, refined by LLM-supervised summarization. Bytedance’s HLLM(bytedanceHLLM) decouples content/behavior analysis via Item/User LLMs, improving sequential modeling accuracy.
In sequential recommendation, generative models—particularly sequence-to-sequence architectures—have emerged as a transformative paradigm by framing next-item prediction as autoregressive generation. Early Transformer-based approaches like SASRec demonstrated the efficacy of self-attention in modeling masked user sequences. Building on this, google(googleTiger) introduces a content-quantized framework encoding item semantics into discrete semantic-ID tuples, with a decoder generating these tokens to enhance cold-start robustness and diversity control. Tencent(tencentLCRec) bridges language and collaborative signals via vector quantization and alignment tuning, better integrating recommendation-specific features into LLMs. Kuaishou(kuaishouOnerec) further unifies retrieval and ranking within a single model through sparse MoE, session-wise generation, and DPO-based alignment. Recent works extend these ideas: Meta(metaHSTU) treats recommendation as sequential transduction, enabling efficient scaling in high-cardinality environments(zhang2023scalinglawlargesequential), while others(thuE4SRec) combines sequential pretraining with LLM instruction tuning by embedding item IDs into prompts, enabling both accurate retrieval and interpretable recommendations.
@@ 3.3. Next Session Prediction Framework
Given TBGRecall’s alignment with mainstream generative models, we integrate two established LLM techniques—Multi-Token Prediction (MTP) and Mixture-of-Experts (MoE)—to enhance performance. We implement Multi-Session Prediction (MSP) along the session dimension for context tokens, mirroring the MTP paradigm (deepseekv3; metamtp). This introduces extended training signals, enabling explicit modeling of long-range user behavior dependencies and transitional relationships across distant contextual scenarios. For the feed-forward module, we adopt DeepSeekMoE with auxiliary-loss-free load balancing, which harnesses specialized expert networks to acquire richer, task-specific knowledge at comparable inference FLOPS, thereby improving overall model efficacy.
@@ 5.1.1. Dataset
We conduct experiments on two large-scale datasets: 1) RecFlow (recflow), which contains rich contextual information from the Kuaishou. 2) TaoBao dataset, a real-world, industrial-scale dataset derived from traffic logs of TaoBao. Table 1 summarizes the main statistics for these two datasets.
@@ 5.1.3. Baselines
To validate the superior performance of our model, we selected multiple generative recommendation baseline models. The models we selected are all non-LLM architectures that, like TBGRecall, utilize sparse IDs. HSTU, SASRec (sasrec), and BERT4Rec (bert4rec) are sequential modeling-based recommendation models, while YouTubeDNN (youtubednn) and FDSA (fdsa) represent traditional recommendation approaches. The ONLINE model adopts Taobao’s production-level dual-tower baseline for recommendation retrieval. The ONLINE (DT) denotes that the model is updated through daily training over multiple years, ensuring sustained adaptation to long-term trends and dynamic user preferences.
@@ 5.3. Ablation Study
We conducted comprehensive ablation experiments on key components including TSN, MSP, MoE(deepseekv3), and RoPE(rope) to assess their individual contributions to model performance. Results confirm that each module independently enhances retrieval accuracy, with the full framework achieving the highest overall effectiveness. MSP contributes the most significant performance improvement, primarily because it increases the amount of information encoded within a single session, enabling the model to better understand the contextual semantics and avoid falling into local decision-making patterns during training. Both session-wise RoPE (Rotary Position Embedding) and MoE (Mixture of Experts) are widely adopted techniques in current state-of-the-art large language models. By incorporating them into our architecture, we achieve considerable gains in model performance. Additionally, the TSN (Token-Specific Network) module enables context tokens and item tokens to align more effectively and efficiently, further enhancing the overall performance of the model. These findings validate the architectural soundness and practical value of the proposed components in advancing generative retrieval systems for e-commerce recommendation.